# Multilingual Prompting for Improving LLM Generation Diversity

### v1

@@ Abstract
Large Language Models (LLMs) are known to lack cultural representation and overall diversity in their generations, from expressing opinions to answering factual questions. To mitigate this problem, we propose multilingual prompting: a prompting method which generates several variations of a base prompt with added cultural and linguistic cues from several cultures, generates responses, and then combines the results. Building on evidence that LLMs have language-specific knowledge, multilingual prompting seeks to increase diversity by activating a broader range of cultural knowledge embedded in model training data. Through experiments across multiple models (GPT-4o, GPT-4o-mini, LLaMA 70B, and LLaMA 8B), we show that multilingual prompting consistently outperforms existing diversity-enhancing techniques such as high-temperature sampling, step-by-step recall, and personas prompting. Further analyses show that the benefits of multilingual prompting vary with language resource level and model size, and that aligning the prompting language with the cultural cues reduces hallucination about culturally-specific information.
@@ 1 Introduction
Figure 1: An example of the diversity of an LLM (GPT-4o)’s responses when prompted in English versus in multiple languages: on the left, we show demographic diversity, specifically the range of different nationalities represented in an answer abut what singers to follow; on the right, we have the level of agreement to a controversial social norms question. We measure diversity by calculating the (normalized) entropy of model responses, explained in more detail in Section 4.1. Multilingual prompting leads to an increase in diversity.
@@
Figure 2: Above: overview of multilingual and multicultural prompting procedure, and our diversity evaluation. Below: example prompts from our multilingual and multicultural methods, and a subset of comparison methods.
@@
Large Language Models (LLMs) are now omnipresent: they have effectively replaced traditional search engines, and people use them to do everything from study to plan their travel and other leisure activities. As a result, LLMs have an ever-increasing power of dictating exposure of ideas, facts, and people as the public use LLMs to gain access to information. It is important that this exposure is distributed in an equitable manner. Lack of diversity in LLM generations—especially when querying for new information—can lead to a host of problems: lack of demographic diversity when queried about individuals can lead to unfair lack of exposure of artists, academics, and other professionals on the basis of their race, ethnicity, or nationality. Lack of cultural diversity when asking opinions on controversial topics can contribute to inaccurate results when using LLMs as substitutes for human responses in user studies, annotation tasks, and opinion surveys, as they do not reflect the diversity of real-world perspectives. Indeed, prior work has shown that LLMs do not represent the true diversity of human expression in a variety of ways—from reducing sentiment and topic diversity for tasks such as book reviews Wu et al. (2024), to demonstrating poor linguistic diversity when helping humans write essays Padmakumar and He (2023). Perhaps even more importantly, LLMs have been shown to generate largely monocultural responses to controversial questions, often leaning towards expressing Western values Wang et al. (2025)—or even a subset of Western values Santurkar et al. (2023).
@@
These trends continue in our own experiments. As we see in Figure 1, when prompted in English, LLMs concentrate their responses about individuals in various professions, e.g., what artists to listen to, to largely American artists, systematically over-looking those from underrepresented cultural backgrounds. Similarly, when we ask LLMs in English whether they agree with a statement known to be controversial among humans Forbes et al. (2020), e.g., “It is ok to live with a roommate of the opposite sex if you are just friends”— the models largely agree with this statement, generating homogeneous responses which do not reflect the variety of perspectives across different cultural contexts.
@@
In this work, we propose that language and other cultural cues can be a powerful lever for enhancing diversity in LLM outputs, and thus point to a way to mitigate these problems. Returning to Figure 1, we see that prompting the model in multiple languages at once and combining the responses, leads to higher diversity in the ethnicity and nationality of the artists suggested. Similarly, if we ask the model in several different languages about living with opposite sex roommates, the response varies much more. These results add to increasing evidence Aggarwal et al. (2025); Hämäläinen et al. (2023) that LLMs encode culturally specific information linked to the language and other cultural cues in the input–—and we suggest these differences in LLM behavior across different languages and cultural cues present an opportunity to deliberately create more diverse generations.
@@
But this begs the question: what is the best way to prompt the model to tap into its culture-specific knowledge, in order to create more diverse, but correct, generations? Is language itself the best signal to prompt the model to dip into particular cultural knowledge, or are cultural cues such as giving a name, birthplace and personality cues for a persona on their own enough? (See Figure 2 for example prompts). In Sections 4 and 6, we explore these questions, and find that both language and cultural cues are important for boosting diversity, but prompting in the language connected to a given culture achieves higher diversity overall, and is important to prevent hallucination for culturally relevant information, e.g. giving the names of actual Chinese singers as opposed to random Chinese names.
@@
In sum, in this work, we present the following three contributions: (1) we introduce and evaluate multilingual and multicultural prompting and shown in 2 as methods to increase various forms of demographic, cultural, and other forms of diversity in LLM generations. We find that these methods increase demographic and cultural diversity in LLM generations better than state of the art methods such as step-by-step recall prompting Hayati et al. (2023), generating personas Wang et al. (2025), and increasing temperature Chung et al. (2023), all while maintaining accuracy on factual tasks. (2) We explore whether using the native language that corresponds to the cultural cues reduces hallucination for culture-specific pieces of information, such as names of famous singers from different parts of the world. Based on human evaluation of model outputs, we find that specifically prompting in the language associated with a specific culture reduces hallucinations about that culture when compared to prompting in English, suggesting that language is imperative for generating accurate and diverse information. (3) Finally, we evaluate the performance of multilingual prompting as the number of languages increases, as well as across lower- and high-resourced languages. We see that overall, the diversity gain from multilingual prompting increases with the number of languages used. Further, we see that some models gain more diversity from prompting in high-resourced languages, while smaller models demonstrate greater diversity gains from lower-resourced languages.
@@ 2 Related Work
Current Diversity Issues in LLMs. Recent research has raised concerns about lack of diversity in LLM opinions, cultural perspective, and linguistic expression Wang et al. (2025); Padmakumar and He (2023); Tevet and Berant (2020). For example, recent work has revealed that LLMs reflect opinions of dominant groups disproportionately even despite prompt steering (Santurkar et al., 2023), and that LLMs can produce nearly identical responses even when primed with demographic variation in prompts (Park et al., 2024; Kitadai et al., 2024). More broadly, many authors have expressed concern about homogenizing, monocultural tendencies of LLMs leading to societal harm, from discrimination to model collapse (Bommasani et al., 2022; Fabris et al., 2022; Kleinberg and Raghavan, 2021; Wu et al., 2024; Shumailov et al., 2024).
@@
To counter these issues, researchers have explored methods to increase diversity in LLM outputs while maintaining coherence and accuracy. We compare multilingual prompting to many of these methods in Section 4.2, including sampling-based approaches (e.g., high temperature, top-$k$ sampling); persona-based prompting (Cheng et al., 2023), where models simulate varied viewpoints by adopting socio-demographic roles or synthetic identities (Mukherjee et al., 2024; Beck et al., 2023); and step-by-step recall prompting, which encourages the model to explore multiple evaluative dimensions or iteratively expand its answer space (Hayati et al., 2023). Overall, based on our evaluation of demographic diversity for prompts about individuals and diversity of perspective in prompts on social norms, we find that multilingual prompting is more effective than these other methods.
@@
LLMs Across Languages. A separate line of work has shown that LLMs perform variably across languages Ohmer et al. (2023); Goldman et al. (2025). While much of this work has focused on negatives—–e.g., showing that LLMs have differing ability to recall facts in different languages–—we argue that this variability can be exploited. Perhaps most related, Kwok et al. explore to what extent language and other cultural cues can help LLMs respond to questions in a manner that reflect a particular cultural background Kwok et al. (2024). Importantly, our work differs in that we suggest multilingual prompting as a method to improve general diversity in LLM responses, rather than attempting to faithfully recreate a particular cultural background. Interestingly, their findings suggest that using native language is not helpful for eliciting representative responses for specific cultures, but that culture-and nationality-specific cues in English are most effective. However, we find that adding native language provides a diversity boost when used in conjunction with cultural cues. Further, while Kwok et al. Kwok et al. (2024) find that using native language decreases performance of matching human outputs from a given culture, we find that using native language increases performance of the LLM by decreasing culture-specific hallucination (see Section 5).
@@ 3 Multilingual and Multicultural Prompting
We present two related prompting methods in this work, which we call multilingual and multicultural prompting. Both multilingual and multicultural prompting work to increase LLM generation diversity by eliciting responses to several different version of the same prompt, each with different cultural and/or linguistic cues, and then combining them into one response. One goal of this work to understand which method is the best to increase diversity in LLM generations. Multicultural prompting does so by relying solely on adding cultural cues, in English—such as adding to the prompt that the LLM is English-speaking, or giving a persona with a Chinese name and adding they were born in Beijing. For multilingual prompting, we rely on these cultural cues and translating the prompt to the language associated with that culture. See Figure 2 for examples. Our multilingual and multicultural prompting methods consist of three main steps, also shown in Figure 2:
@@
1. Preparation of Queries: We begin by editing the original English query by creating $n$ versions of the original query, each with added cues related to various languages or culture (e.g., "You are a Chinese-speaking assistant", see more in Figure 2) and, in the case of multilingual prompting, also translating the prompt into the corresponding target languages (e.g. Chinese). For example, to do multilingual or multicultural prompting with English, Chinese, and Japanese, we generate three versions of the prompt, each corresponding to one language and cultural background.
@@
We have two types of multicultural/lingual queries, one set of which we label “basic” and the other we label“enhanced”. The basic consists of only the cultural cue “[language]-speaking. The enhanced multicultural/lingual prompting comes from adding three addition cultural cues: a name, birthplace, and favorite food. Importantly, part of even our basic multilingual prompting technique includes a cue that the model speaks the language in question, in our case, "You are an [language]-speaking assistant". Following prior work Kwok et al. (2024), in preliminary experiments we find that language completely on its own does not increase diversity.
In our open-source application [^1] Code will be released pending acceptance., users can select arbitrary target languages to suit their own cultural preferences. For demonstration purposes, we have chosen Chinese, Japanese, and English in our experiments, as the authors speak all three languages.
@@
2. Model Response Generation: The modified prompts (one per language) are then given to the LLM one at a time. The model generates responses for each modified query. For multilingual prompting, the model responds in various different languages, and we translate all answers back into English using GPT-4o-mini.
3. Aggregation: We then combine the responses into one answer. In this work, we concatenate the responses to most easily tabulate diversity and compare multilingual prompting with other methods. However, more broadly, we suggest three methods for combining the prompts depending on their use case: concatenation, summarization, or random selection. Concatenation is simply adding the responses together, summarization involves asking the model to summarize the three responses in English, and random selection involves selecting from the three answers with some distribution of probability. While we present results on concatenation in this work, we discuss these methods in more detail in the discussion and limitations section.
@@ 4 Increasing Demographic and Perspective Diversity
In this section, we present our experimental framework and results showing how multilingual prompting can increase demographic and perspective diversity compared to other state of the art methods.
4.1 Experimental Setup
@@
To evaluate LLM response diversity across various prompting strategies, we test the diversity of LLM responses for two tasks: responding to questions about controversial social norms, and naming individuals in various professions.
@@
Social Norms: Social-Chem 101 Dataset. For this task, we ask the LLM to select a multiple-choice response to a statement making a claim about a controversial social norm—i.e., one that humans have differing opinions on. The multiple choice answers encode agreement or disagreement, as well as reasoning. We evaluate diversity for each of these, as described below.
@@
Specifically, we use a modified version of that the Social Chemistry 101 dataset Forbes et al. (2020), a large-scale corpus designed to capture human moral and social norms through Rules of Thumb (RoTs)—natural language statements reflecting implicit principles of social acceptability. While original dataset only includes statements themselves, in order to evaluate diversity in model agreement and model reasoning, we used an LLM to generate multiple choice responses to the statements. We generated four choices, two agrees and two disagrees, and explanations for each questions that encoded different moral values. Examples of these answers, and the prompts used to generate them, are in Appendix A.1.1. We then use these multiple choice answers to calculate both the diversity in agreement and in the reason behind the judgment. Example statements are “You should not feel obligated to go to your family reunions” and “It’s wrong to keep your health a secret”.
The original dataset includes human baseline judgments, where annotators indicate the degree of agreement with each statement. In this study, in order to ensure we are focusing on truly controversial statements where diverse responses are desirable, we specifically focus on a subset of the dataset where between only 25-75% of annotators agree with the given statement.
@@
Demographic Representation: People Diversity Dataset. In this task, we ask an LLM to generate names for individuals in various professions, and assess the demographic diversity of the response across nationality, ethnicity, and geographic region. To do so, we use the People Diversity Dataset Lahoti et al. (2023). The dataset consists of 105 questions asking for names of individuals across different professions in a context, for example, “Name some writers that inspire you”. In our experiment, we ask the LLMs to generate exactly 10 names for each question.
@@
Metrics. To assess the diversity of LLM responses to questions from both datasets, we rely on different calculations of entropy.
@@
Reason and Valence Entropy. To assess the diversity in reasoning across LLM responses to social norms questions, for each prompting strategy, we calculate the average entropy across the three responses we generate from the model for each query corresponding to each language or culture. We call this reason entropy. To assess the diversity in agreement/disagreement, we calculate entropy but treat responses that have the same decision (agree/disagree) as interchangeable. We call this valence entropy. For example, Reason Entropy is calculated as $H_{\text{Reason}}=-\sum_{i\in\{A,B,C,D\}}p(i)\log p(i)$, where $p(i)$ represents the probability of the model selecting choice $i$. Valence entropy only has two choices, agree or disagree. A higher entropy indicates a greater diversity.
@@
Demographic Entropies. To evaluate demographic representation, we use an LLM (GPT-4o-mini) to annotate the nationality, ethnicity, and region for each name generated. To ensure the reliability of these cultural origin annotations, we conduct performance checks for the annotation (See details in Appendix A.2.2). Then, for each question, we calculate the entropy for each attribute across the thirty names generated from each prompting strategy to measure the cultural diversity of the model’s predictions. For each attribute $A$, we define its entropy $H(A)$ as: $H(A)=-\sum_{c\in C_{A}}p(c)\log p(c)$, where $C_{A}$ is the set of possible categories within attribute $A$, and $p(c)$ represents the probability of category $c$ occurring in the model’s annotations. In Section 4.2, we report the average normalized entropy across all questions.
@@
To place all metrics on a common $[0,1]$ scale, we divide each raw score by the maximum value it could theoretically attain under the same option count, $\widetilde{H}\;=\;\frac{H}{H_{\max}}$, where $H$ is the unnormalized value and $H_{\max}$ is the corresponding upper bound. More details about normalization and metrics for further experiments can be found in Appendix A.3 and Appendix A.4.

### v2

@@ Abstract
Large Language Models (LLMs) are known to lack cultural representation and overall diversity in their generations, from expressing opinions to answering factual questions. To mitigate this problem, we propose multilingual prompting: a prompting method which generates several variations of a base prompt with added cultural and linguistic cues from several cultures, generates responses, and then combines the results. Building on evidence that LLMs have language-specific knowledge, multilingual prompting seeks to increase diversity by activating a broader range of cultural knowledge embedded in model training data. Through experiments across multiple models (GPT-4o, GPT-4o-mini, LLaMA 70B, and LLaMA 8B), we show that multilingual prompting consistently outperforms existing diversity-enhancing techniques such as high-temperature sampling, step-by-step recall, and persona prompting. Further analyses show that the benefits of multilingual prompting vary between high and low resource languages and across model sizes, and that aligning the prompting language with cultural cues reduces hallucination about culturally-specific information.
@@ 1 Introduction
Figure 1: An example of the diversity of an LLM’s (GPT-4o) responses when prompted in English versus in multiple languages: on the left, we show demographic diversity, specifically the range of different nationalities represented in an answer about which singers to follow; on the right, we show the level of agreement with a controversial social norms statement. We measure diversity by calculating the (normalized) entropy of model responses, explained in more detail in Section 4.1. Multilingual prompting leads to an increase in diversity.
@@
Figure 2: Above: an overview of multilingual and multicultural prompting, and our diversity evaluation. Below: example prompts from our multilingual and multicultural methods, and a subset of methods we compare to.
@@
Large Language Models (LLMs) are now omnipresent: they have effectively replaced traditional search engines, and people use them to do everything from studying to planning travel and other leisure activities Chatterji et al. (2025). As a result, LLMs have an ever-increasing power to dictate exposure to ideas, facts, and people, as the public uses LLMs to gain access to information. It is important that this exposure is distributed in an equitable manner. Lack of diversity in LLM generations—especially when querying for new information—can lead to a host of problems: lack of demographic diversity when the LLM is queried about individuals can lead to unfair lack of exposure of artists, academics, and other professionals on the basis of their race, ethnicity, or nationality. Lack of cultural diversity in response to questions about controversial topics can contribute to inaccurate results when LLMs are used as substitutes for human responses in user studies, annotation tasks, and opinion surveys, as these responses do not reflect the diversity of real-world perspectives. Indeed, prior work has shown that LLMs do not represent the true diversity of human expression in a variety of ways—from reducing sentiment and topic diversity for tasks such as book reviews Wu et al. (2024), to demonstrating poor linguistic diversity when helping humans write essays Padmakumar and He (2023). Perhaps even more importantly, LLMs have been shown to generate largely monocultural responses to controversial questions, often leaning toward expressing Western values Wang et al. (2025), or even a subset of Western values Santurkar et al. (2023).
@@
These trends continue in our own experiments. In Figure 1, we show an example of LLM responses when asked about individuals in various professions, for example, what musical artists to listen to. When prompted in English, LLM answers are largely limited to American artists, and exclude those from other cultural backgrounds. Similarly, when we ask LLMs in English whether they agree with a statement known to be controversial among humans Forbes et al. (2020), e.g., “You should not feel obligated to go to your family reunions”— the models largely agree with this statement, generating homogeneous responses which do not reflect the variety of perspectives across different cultural contexts.
@@
In this work, we propose that language and other cultural cues can be a powerful lever for enhancing diversity in LLM outputs, which points to a way to mitigate these problems. Returning to Figure 1, we see that prompting the model separately in multiple different languages and combining the responses leads to higher diversity in the ethnicity and nationality of the artists suggested. Similarly, if we ask the model in several different languages for opinions about whether people should feel obligated to attend their family reunions, the response varies much more in its level of agreement with the statement. These results add to increasing evidence Aggarwal et al. (2025); Hämäläinen et al. (2023) that LLMs encode culturally specific information linked to the language and other cultural cues in the input–—and we suggest these differences in LLM behavior across different languages and cultural cues present an opportunity to deliberately create more diverse generations.
@@
But this raises the question: what is the best way to prompt the model to tap into its culture-specific knowledge, in order to create more diverse, but correct, generations? Is language itself the best signal to prompt the model to dip into particular cultural knowledge, or are cultural cues such as giving a name, birthplace and personality cues for a persona on their own enough? (See Figure 2 for example prompts.) In Sections 4 and 6, we explore these questions. We find that both language and cultural cues are important for boosting diversity, but prompting in the language connected to a given culture achieves higher diversity overall. Further, we see that matching cultural cues and language is important to prevent hallucination for culturally relevant information, e.g. giving the names of actual Chinese singers as opposed to simply outputting Chinese names.
@@
In sum, in this work, we present the following three contributions: (1) we introduce and evaluate multilingual and multicultural prompting as shown in Figure 2 as methods to increase demographic, cultural, and other forms of diversity in LLM generations. We find that these methods increase demographic and cultural diversity in LLM generations better than state of the art methods, such as step-by-step recall prompting Hayati et al. (2023), generating personas Wang et al. (2025), and increasing temperature Chung et al. (2023), all while maintaining accuracy on factual tasks. (2) We explore whether using the native language that corresponds to the cultural cues reduces hallucination for culture-specific pieces of information, such as names of famous singers from different parts of the world. Based on human evaluation of model outputs, we find that specifically prompting in the language associated with a specific culture reduces hallucinations about that culture when compared to prompting in English, suggesting that language is imperative for generating accurate and diverse information. (3) Finally, we evaluate the performance of multilingual prompting as the number of languages increases, as well as across lower- and high-resourced languages. We see that overall, the diversity gain from multilingual prompting increases with the number of languages used. Further, we see that some models gain more diversity from prompting in high-resourced languages, while smaller models demonstrate greater diversity gains from lower-resourced languages.
@@ 2 Related Work
Current Diversity Issues in LLMs. Recent research has raised concerns about the lack of diversity in LLM opinions, cultural perspectives, and linguistic expression Wang et al. (2025); Padmakumar and He (2023); Tevet and Berant (2020). For example, recent work has revealed that LLMs reflect the opinions of dominant groups disproportionately even despite prompt steering (Santurkar et al., 2023), and that LLMs can produce nearly identical responses even when primed with demographic variation in prompts (Park et al., 2024; Kitadai et al., 2024). More broadly, many authors have expressed concern about homogenizing, monocultural tendencies of LLMs leading to societal harm, from discrimination to model collapse (Bommasani et al., 2022; Fabris et al., 2022; Kleinberg and Raghavan, 2021; Wu et al., 2024; Shumailov et al., 2024; Zhang et al., 2024).
@@
To counter these issues, researchers have explored methods to increase diversity in LLM outputs while maintaining coherence and accuracy. We compare multilingual prompting to many of these methods in Section 1, including sampling-based approaches (e.g., high temperature, top-$k$ sampling); persona-based prompting (Cheng et al., 2023), where models simulate varied viewpoints by adopting socio-demographic roles or synthetic identities (Mukherjee et al., 2024; Beck et al., 2023); and step-by-step recall prompting, which encourages the model to explore multiple evaluative dimensions or iteratively expand its answer space (Hayati et al., 2023). Overall, based on our evaluation of demographic diversity for prompts about individuals and diversity of perspective in prompts on social norms, we find that multilingual prompting is more effective than these other methods.
@@
LLMs Across Languages. A separate line of work has shown that LLMs perform variably across languages Ohmer et al. (2023); Goldman et al. (2025). While much of this work has focused on negatives—–e.g., showing that LLMs have differing ability to recall facts in different languages—we argue that this variability can be exploited. Perhaps most related to our work, Kwok et al. explore to what extent language and other cultural cues can help LLMs respond to questions in a manner that reflect a particular cultural background Kwok et al. (2024). Importantly, our work differs in that we suggest multilingual prompting as a method to improve general diversity in LLM responses, rather than attempting to faithfully recreate a particular cultural background. Interestingly, their findings suggest that using native language is not helpful for eliciting representative responses for specific cultures, but that culture-and nationality-specific cues in English are most effective. However, we find that adding native language provides a diversity boost when used in conjunction with cultural cues. Further, while Kwok et al. (2024) find that using native language decreases performance of matching human outputs from a given culture, we find that using native language increases the performance of the LLM by decreasing culture-specific hallucination (see Section 5).
@@ 3 Multilingual and Multicultural Prompting
We present two related prompting methods in this work, which we call multilingual and multicultural prompting. Both multilingual and multicultural prompting work to increase LLM generation diversity by eliciting responses to several different versions of the same prompt, each with different cultural and/or linguistic cues, and then combining them into one response. One goal of this work is to understand which method is the best to increase diversity in LLM generations. Multicultural prompting does so by relying solely on adding cultural cues, in English—such as adding to the prompt that the LLM is English-speaking, or giving a persona a Chinese name and adding they were born in Beijing. For multilingual prompting, we rely on these cultural cues and translate the prompt to the language associated with that culture. See Figure 2 for examples. Our multilingual and multicultural prompting methods consist of three main steps, also shown in Figure 2:
1. Preparing the Queries:
@@
We begin by editing the original English query by creating $n$ versions of the original query, each with added cues related to various languages or culture (e.g., “You are a Chinese-speaking assistant”, see more in Figure 2) and, in the case of multilingual prompting, also translating the prompt into the corresponding target languages (e.g. Chinese). For example, to do multilingual or multicultural prompting with English, Chinese, and Japanese, we generate three versions of the prompt, each corresponding to one language and set of cultural cues.
@@
We have two types of multicultural/lingual queries, one set of which we label “basic” and the other we label “enhanced”. The basic variant consists of prompting the model with “You are an [language]-speaking assistant”. Enhanced multicultural/lingual prompting adds three addition cultural cues: a name, birthplace, and favorite food. Following prior work Kwok et al. (2024), in preliminary experiments, we find that language completely on its own, without any cultural signal, does not increase diversity.
In the implementation we release, [^1] Our code repository is available at: https://github.com/mangocyann/Multilingual-Prompting-for-Improving-LLM-Generation-Diversity. users can select arbitrary target languages to suit their own cultural preferences. The experiments presented in this paper focus on Chinese, Japanese, and English in our initial experiments in Sections 4, before expanding to Spanish, French, Nepali, Thai, Turkish, and Ukrainian in section 6.
2. Model Response Generation:
@@
The modified prompts (one per language) are then given to the LLM one at a time. The model generates responses for each modified query. For multilingual prompting, the model responds in various different languages, and we translate all answers back into English using GPT-4o-mini.
3. Aggregation:
We then combine the responses into one answer. In this work, we simply concatenate the responses from the modified prompts to generate one overall response. We make this choice easily by tabulating diversity by comparing the range of responses from multilingual prompting to other prompting methods, such as our baseline of rephrasings of the original query in English. We discuss other methods for aggregation of LLM responses in our discussion section.
@@ 4 Increasing Demographic and Perspective Diversity
In this section, we present our experimental framework and results showing how multilingual prompting can increase demographic and perspective diversity compared to other methods.
4.1 Dataset and Metrics
@@
We test the diversity of LLM responses for two tasks: responding to questions about controversial social norms, and naming individuals in various professions.
@@
Social Norms: the Social Chemistry 101 Dataset. For this task, we ask the LLM to select a multiple-choice response to a statement making a claim about a controversial social norm—i.e., one that humans have differing opinions on.
@@
Specifically, we use a modified version of the Social Chemistry 101 dataset Forbes et al. (2020), a large-scale corpus designed to capture human moral and social norms through Rules of Thumb—natural language statements reflecting implicit principles of social acceptability. The dataset includes human baseline judgments, where annotators indicate the degree of agreement with each statement. In order to ensure we are focusing on truly controversial statements where diverse responses are desirable, we specifically focus on a subset of the dataset where only 25–75% of annotators agree with the given statement. Example statements include “You should not feel obligated to go to your family reunions” and “It’s wrong to keep your health a secret”.
While the original dataset only includes the statements themselves, in order to evaluate diversity in model agreement and model reasoning, we used an LLM to generate multiple-choice responses to each statement. Specifically, we generated four choices, including two agreeing and two disagreeing valences, all with explanations that encoded different moral values to support the viewpoints. For example, “valence”: “agree”; “explanation”: “Obligations shouldn’t dictate personal choices; you should only go if you want to.” More examples and the prompts used to generate them, are in Appendix A.1.1. We then use these multiple-choice answers to calculate the diversity in both the judgment (agreeing or disagreeing) and in the justification for it.
@@
Demographic Representation: the People Diversity Dataset. In this task, we ask an LLM to generate names for individuals in various professions, and assess the demographic diversity of the response in terms of nationality, ethnicity, and geographic region. To do so, we use the People Diversity Dataset Lahoti et al. (2023). The dataset consists of 105 questions asking for the names of individuals in a particular profession and context. For example, “Name some writers that inspire you”. In our experiment, we ask the LLMs to generate exactly 10 names for each question.
Metrics.
@@
To assess the diversity of LLM responses to questions from both datasets, we rely on the following entropy measures.
@@
Reason and Valence Entropy. To assess the diversity in reasoning across LLM responses to social norms questions, for each prompting strategy, we calculate the average entropy across the three responses we generate from the model for each query corresponding to each language or culture. We call this reason entropy. To assess the diversity in agreement/disagreement, we calculate entropy, but treat responses that have the same judgment (agree/disagree) as interchangeable. We call this valence entropy. For example, Reason Entropy is calculated as $H_{\text{Reason}}=-\sum_{i\in\{A,B,C,D\}}p(i)\log p(i)$, where $p(i)$ represents the probability of the model selecting choice $i$. Valence entropy only has two choices: agree or disagree. A higher entropy indicates a greater diversity, because we focus on controversial questions, higher diversity is generally desirable.
Demographic Entropies.
@@
To evaluate demographic representation, we use an LLM (GPT-4o-mini) to annotate the nationality, ethnicity, and region for each name generated. To ensure the reliability of these cultural origin annotations, we conduct performance checks for the annotation (see details in Appendix A.2.2). Then, for each question, we calculate the entropy for each attribute across the thirty names (10 names for each of the 3 languages) generated from each prompting strategy to measure the cultural diversity of the model’s predictions. For each attribute $A$, we define its entropy $H(A)$ as: $H(A)=-\sum_{c\in C_{A}}p(c)\log p(c)$, where $C_{A}$ is the set of possible categories within attribute $A$, and $p(c)$ represents the probability of category $c$ occurring in the model’s annotations. In Section 1, we report the average normalized entropy across all questions.
@@
To place all metrics on a common $[0,1]$ scale, we divide each raw score by the maximum value it could theoretically attain under the same option count, $\widetilde{H}\;=\;\frac{H}{H_{\max}}$, where $H$ is the unnormalized value and $H_{\max}$ is the corresponding upper bound. Details about normalization can be found in Appendix A.3 and Appendix A.4.
4.2 Baselines and Performance Tests